OpenStudy (anonymous):

Find two consecutive odd integers such that the sum of their squares is 130.

OpenStudy (anonymous):

n^2 + (n+2)^2 = 130

OpenStudy (anonymous):

2n^2 + n -126 = 0 n^2 + 2n -63=0 now solve it

OpenStudy (anonymous):

the ans is 7 and 9

OpenStudy (anonymous):

thank you umm 3 and 5 or -3 and -5?